OpenStudy (anonymous):

A toy rocket is launched vertically from ground level (y = 0 m), at time t = 0.0 s. The rocket engine provides constant upward acceleration during the burn phase. At the instant of engine burnout, the rocket has risen to 68 m and acquired a velocity of 90 m/s. The rocket continues to rise in unpowered flight, reaches maximum height, and falls back to the ground. The time interval, during which the rocket is in unpowered flight, is closest to: 14 s 15 s 19 s 16 s 18 s

OpenStudy (anonymous):

you have to calculate two time interval. first t1 = time rocket takes to reach at its maximum height from 68m. and second time takes by rocket to free fall from its maximum height to ground add t1 and t2 you will get the answer.

OpenStudy (anonymous):

@gyanu thanks for your responding, but I am still confused a bit, could show me the equation that I can use !

OpenStudy (anonymous):

ok first we get t1 by using equation v=u-gt1 in this equation you have to put v=0 and u=90 m/s now from eqn \[v ^{2}=u ^{2}-2gh\] u can find h for same u and v value given above. now by adding u can get total height H=68+h is distance rocket travel i free fall. u have to use now eqn \[H=ut+\frac{ 1 }{ 2 } g t _{2}^{2}\] by putting u=0 and H=68+h u will find t2. its a lengthy calculation so do calculation care fully.

OpenStudy (anonymous):

Ohh thanks I got the answer which is : 19s